OpenStudy (anonymous):

An ant travels 2.78cm [W] and then turns and travels 6.25cm [S 40 degrees E]. What is the ant's total displacement?

OpenStudy (anonymous):

 You are going to solve for R using the cosine theorem c^2=a^2 + b^2 -2abcos(theta) c^2=(2.78)^2 + (6.25)^2 -2(2.78)(6.25)cos(40) c^2=7.73+39.06- 26.62 c^2=20.17 sqrt(c)= sqrt(20.17) c=4.49cm

OpenStudy (anonymous):

like this,,,